1. Factory Overhead Controllable Variance

Bellingham Company produced 5,400 units of product that required 5.5 standard direct labor hours per unit. The standard variable overhead cost per unit is $6.00 per direct labor hour. The actual variable factory overhead was $170,360. Determine the variable factory overhead controllable variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

2. Factory Overhead Volume Variance

Bellingham Company produced 5,100 units of product that required 1.5 standard direct labor hours per unit. The standard fixed overhead cost per unit is $2.20 per direct labor hour at 7,050 hours, which is 100% of normal capacity. Determine the fixed factory overhead volume variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.
Answer #1

Variable factory overhead controllable variance

= Standard voh cost - Actual voh cost

= (5400*5.5*6) - 170360

= 178200 - 170360

= 7840 favorable,

Fixed factory overhead volume variance

= (Budgeted output - Actual output) * FOAR

= (7050/1.5 - 5100) * (1.5*2.20)

= 1320 Favourable